OpenStudy (anonymous):

Find the height of a rectangular prism if the surface area is 148 cm2, the width is 4 cm and the length is 5 cm. i dont get it

OpenStudy (anonymous):

due to the nature of a rectangular prism each end is a square. refer to my drawing. so area due to squares at either end is 2*4*4=32. That means that the recatangles on the side have a total area of 148-32=116. There are four rectangles so area of each rectangle is 116/4=29. we know the length if the rectangles is 5, as area of rectangle is base* height, 29=5*h, therefore h=5.8cm